If 5 times the smaller of two numbers is subtracted from twice the larger, the result is 16. If the larger is increased by three times the smaller, the result is 63. Find the larger number.
Answer by blwinbbbles(34)

write down what you have..use S for smaller and L for larger
you know that 5S subtracted from 2L is equal to 16
So: 2L - 5S = 16
and you know that L + 3S = 63 and you can write this so L = 63-3S
use substitution:
2(63-3S) - 5S = 16....solve for S
Once you have S then plug it into one of the equations and solve for L
Good Luck